# Analog Computing for Signal Processing and Communications – Part II: Toward Gigantic MIMO Beamforming

### v2

@@ V-C Hybrid Beamforming
In hybrid beamforming [5, 6, 7, 8, 9], the transmitted signal is $\mathbf{x}=\mathbf{W}_{A}\mathbf{W}_{D}\mathbf{s}$, where $\mathbf{W}_{A}\in\mathbb{C}^{N_{T}\times N_{RF}}$ is the analog precoder, subject to specific constraints depending on the architecture of phase shifters and switches, and $\mathbf{W}_{D}\in\mathbb{C}^{N_{RF}\times N_{S}}$ is the digital precoder, subject only to the power constraint, as in Fig. 4(c). Thus, the effective precoding matrix is given by $\mathbf{W}=\mathbf{W}_{A}\mathbf{W}_{D}$. Numerous hybrid beamforming architectures have been proposed, each with different constraints on the matrix $\mathbf{W}_{A}$. For example, in the fully-connected architecture, each RF chain is connected to all the antennas through a phase shifter, and each entry of $\mathbf{W}_{A}$ is subject to the unit-modulus constraint, i.e., $\mathbf{W}_{A}\in\mathbb{T}^{N_{T}\times N_{RF}}$. The computational complexity of hybrid beamforming is given by the complexity of computing the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ on a per-symbol basis, and optimizing $\mathbf{W}_{A}$ and $\mathbf{W}_{D}$ at each coherence block.
@@ V-D RIS-Aided Beamforming
More recently, RIS-aided beamforming emerged by exploiting a RIS deployed in close proximity with an active transmitting device [12, 13, 14, 18], as in Fig. 4(d). The RIS can work in either reflective or transmissive mode, without changing the system model. Denoting as $\mathbf{x}^{\prime}=\mathbf{W}_{D}\mathbf{s}$ the digitally precoded signal transmitted by the active device, the received signal is given by $\mathbf{y}=\mathbf{H}_{\text{EFF}}\mathbf{x}^{\prime}$, where $\mathbf{H}_{\text{EFF}}\in\mathbb{C}^{N_{R}\times N_{RF}}$ is the RIS-aided channel, given by $\mathbf{H}_{\text{EFF}}=\mathbf{H}\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}$, with $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$, $\boldsymbol{\Theta}\in\mathbb{C}^{N_{T}\times N_{T}}$, and $\mathbf{H}_{\text{RIS}}\in\mathbb{C}^{N_{T}\times N_{RF}}$ being the channel from the RIS to the receiver, the scattering matrix of the RIS, and the channel from the active transmitting device to the RIS, respectively. Thus, effective precoding matrix is $\mathbf{W}=\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}\mathbf{W}_{D}$, which is optimized by reconfiguring $\boldsymbol{\Theta}$, subject to $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}_{N_{T}}$ in the case of a lossless RIS [15], and $\mathbf{W}_{D}$, subject only to the power constraint. As in hybrid beamforming, the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ needs to be computed on a per-symbol basis, while $\mathbf{W}_{D}$ and $\boldsymbol{\Theta}$ are optimized per coherence block.
@@ V-E SIM-Aided Beamforming
where $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$, $\boldsymbol{\Theta}_{\ell}\in\mathbb{C}^{N_{T}\times N_{T}}$, and $\mathbf{H}_{\text{SIM},\ell}\in\mathbb{C}^{N_{T}\times N_{T}}$ are the channel from the $L$th RIS to the receiver, the scattering matrix of the $\ell$th RIS, and the channel from the $\ell-1$th to the $\ell$th RIS (where the 00th RIS is the transmitting device and $\mathbf{H}_{\text{SIM},1}\in\mathbb{C}^{N_{T}\times N_{RF}}$), respectively. Thus, the effective precoding matrix reads as
@@ V-F3 Low-resolution ADCs/DACs
Since MiLAC-aided beamforming processes the symbols purely in the analog domain, each RF chain carries an individual symbol, which lies in a constellation with finite cardinality. Thus, low-resolution ADCs/DACs can be adopted without sacrificing performance, which are less expensive and less power-hungry than high-resolution ADCs/DACs. High-resolution ADCs/DACs are instead necessary when beamforming is performed in the digital domain (entirely or in part), i.e., in digital beamforming, hybrid beamforming, and RIS-aided beamforming, as indicated in Tab. IV.
@@ V-F4 No Computations on a Per-Symbol Basis
If beamforming is performed in the digital domain (entirely or in part), the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ needs to be computed on a per-symbol basis, as shown in Tab. IV. The complexity of this operation is $8N_{RF}N_{S}$ real operations, as discussed in the Appendix of Part I of this paper, and can become prohibitive in massive/gigantic MIMO systems [3]. MiLAC-aided beamforming solves this issue since it does not require any computation on a per-symbol basis, as the beamforming is purely analog. This reduces the beamforming computational complexity, especially when the coherence time is much longer than the symbol time [3].
@@ VI-A Performance Evaluation
To evaluate the performance of R-ZFBF, ZFBF, and MBF realized with digital and MiLAC-aided beamforming, we consider a multi-user multiple-input single-output (MISO) system with an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna receivers, as presented in Section IV-B. Here, the transmitter precodes the $N_{R}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{R}\times 1}$ intended to the $N_{R}$ receivers through R-ZFBF or ZFBF to suppress inter-stream interference, or through MBF to maximize the channel gains. With digital beamforming, the R-ZFBF transmitter is realized by computing $\mathbf{F}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$ and setting the precoding matrix $\mathbf{W}_{\text{R-ZFBF}}$ such that $[\mathbf{W}_{\text{R-ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, assuming uniform power allocation to the users. For large $N_{R}$ and if the signal-to-noise ratio (SNR) is the same at all receiving antennas, the optimal $\lambda$ is $\lambda=N_{R}\sigma_{\mathbf{n}}^{2}/P_{T}$, where $P_{T}$ is the transmit power, i.e., $\mathbf{s}$ has covariance matrix $\mathbf{C}_{\mathbf{s}}=P_{T}/N_{R}\mathbf{I}_{N_{R}}$, $\sigma_{\mathbf{n}}^{2}$ is the noise power, and we assumed the channel to have unit path gain [35, Chapter 12]. Similarly, digital beamforming can realize ZFBF by computing $\mathbf{F}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$ and setting $\mathbf{W}_{\text{ZFBF}}$ as $[\mathbf{W}_{\text{ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, and MBF by using $\mathbf{F}_{\text{MBF}}=\mathbf{H}^{H}$ and $\mathbf{W}_{\text{MBF}}$ such that $[\mathbf{W}_{\text{MBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{MBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{MBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$. When a MiLAC is used to perform R-ZFBF, ZFBF, and MBF, we have two choices. First, the MiLAC can be reconfigured to implement an arbitrary beamforming matrix, as discussed in Section IV-A. Thus, its tunable admittance components can be set according to (19) and (20), where $\mathbf{W}$ is the desired beamforming matrix $\mathbf{W}_{\text{R-ZFBF}}$, $\mathbf{W}_{\text{ZFBF}}$, or $\mathbf{W}_{\text{MBF}}$ with uniform power allocation, which is digitally computed. Second, the MiLAC can directly compute the precoding matrices of these LMMSE-inspired beamforming techniques in the analog domain, as explained in Section IV-B. In this case, the columns of $\mathbf{W}_{\text{R-ZFBF}}$, $\mathbf{W}_{\text{ZFBF}}$, and $\mathbf{W}_{\text{MBF}}$ cannot be individually normalized, and the resulting precoding matrices are given by $\mathbf{W}_{\text{R-ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{R-ZFBF}}/\|\mathbf{F}_{\text{R-ZFBF}}\|_{F}$, $\mathbf{W}_{\text{ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{ZFBF}}/\|\mathbf{F}_{\text{ZFBF}}\|_{F}$, and $\mathbf{W}_{\text{MBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{MBF}}/\|\mathbf{F}_{\text{MBF}}\|_{F}$, where the normalization factor has been included to ensure that the transmit power is the same as with digital beamforming.
@@
In Fig. 5, we report the sum rate versus the SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$ achieved by digital and MiLAC-aided beamforming, where $N_{T}=N_{R}=4$ and the channels are independent and identically distributed (i.i.d.) Rayleigh distributed with unit path gain. We observe that R-ZFBF is the best precoder, which boils down to ZFBF at high SNR and to MBF at low SNR, for both digital and MiLAC-aided beamforming, as expected from MIMO theory [35, Chapter 12]. MiLAC-aided beamforming performs exactly as digital beamforming when the MiLAC is reconfigured based on the arbitrary beamforming matrix computed digitally, denoted as MiLAC (arbit.) in Fig. 5. Besides, MiLAC-aided beamforming suffers a slight performance degradation at high SNR when the LMMSE-inspired beamforming matrix is computed in the analog domain, referred to as MiLAC (LMMSE) in Fig. 5. This occurs because, in this case, MiLAC-aided beamforming cannot perform uniform power allocation, which is optimal at high SNR. At the cost of this slight performance degradation, computing the LMMSE-inspired beamforming matrices in the analog domain substantially reduces the required computational complexity, as it will be analyzed in the following.
@@
In Fig. 6, we report the BER versus the SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$ obtained with digital and MiLAC-aided beamforming with $N_{T}=N_{R}=4$ and quadrature phase shift keying (QPSK) modulation. Since MiLAC-aided beamforming can implement the same combining matrices as digital beamforming, the two beamforming strategies achieve exactly the same BER.
@@
Figure 8: Sum rate versus computational complexity of R-ZFBF performed in a $N_{R}\times N_{T}$ system with digital and MiLAC-aided beamforming, for three different levels of SNR.

### v3

@@ V-C Hybrid Beamforming
In hybrid beamforming [6, 7, 8, 9, 10], the transmitted signal is $\mathbf{x}=\mathbf{W}_{A}\mathbf{W}_{D}\mathbf{s}$, where $\mathbf{W}_{A}\in\mathbb{C}^{N_{T}\times N_{RF}}$ is the analog precoder, subject to specific constraints depending on the architecture of phase shifters and switches, and $\mathbf{W}_{D}\in\mathbb{C}^{N_{RF}\times N_{S}}$ is the digital precoder, subject only to the power constraint, as in Fig. 4(c). Thus, the effective precoding matrix is given by $\mathbf{W}=\mathbf{W}_{A}\mathbf{W}_{D}$. Numerous hybrid beamforming architectures have been proposed, each with different constraints on the matrix $\mathbf{W}_{A}$. For example, in the fully-connected architecture, each RF chain is connected to all the antennas through a phase shifter, and each entry of $\mathbf{W}_{A}$ is subject to the unit-modulus constraint, i.e., $\mathbf{W}_{A}\in\mathbb{T}^{N_{T}\times N_{RF}}$. The computational complexity of hybrid beamforming is given by the complexity of computing the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ on a per-symbol basis, and optimizing $\mathbf{W}_{A}$ and $\mathbf{W}_{D}$ at each coherence block. This is because the digitally precoded signal $\mathbf{W}_{D}\mathbf{s}$ is multiplied by $\mathbf{W}_{A}$ in the analog domain. Although the network of phase shifters in hybrid beamforming can operate a matrix-vector product in the analog domain, it cannot compute non-linear operations as a MiLAC made of admittance components.
@@ V-D RIS-Aided Beamforming
As a related analog-domain beamforming strategy, RIS-aided beamforming emerged by exploiting a RIS deployed in close proximity with an active transmitting device [14, 15, 16, 20], as in Fig. 4(d). The RIS can work in either reflective or transmissive mode, without changing the system model. Denoting as $\mathbf{x}^{\prime}=\mathbf{W}_{D}\mathbf{s}$ the digitally precoded signal transmitted by the active device, the received signal is given by $\mathbf{y}=\mathbf{H}_{\text{EFF}}\mathbf{x}^{\prime}$, where $\mathbf{H}_{\text{EFF}}\in\mathbb{C}^{N_{R}\times N_{RF}}$ is the RIS-aided channel, given by $\mathbf{H}_{\text{EFF}}=\mathbf{H}\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}$, with $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$, $\boldsymbol{\Theta}\in\mathbb{C}^{N_{T}\times N_{T}}$, and $\mathbf{H}_{\text{RIS}}\in\mathbb{C}^{N_{T}\times N_{RF}}$ being the channel from the RIS to the receiver, the scattering matrix of the RIS, and the channel from the active transmitting device to the RIS, respectively. Thus, effective precoding matrix is $\mathbf{W}=\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}\mathbf{W}_{D}$, which is optimized by reconfiguring $\boldsymbol{\Theta}$, subject to $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}_{N_{T}}$ in the case of an ideal lossless RIS [17], $\mathbf{H}_{\text{RIS}}$ is not reconfigurable, and $\mathbf{W}_{D}$, subject only to the power constraint. As in hybrid beamforming, the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ needs to be computed on a per-symbol basis, while $\mathbf{W}_{D}$ and $\boldsymbol{\Theta}$ are optimized per coherence block.
@@ V-E SIM-Aided Beamforming
[…]
@@ V-F3 Low-resolution ADCs/DACs
Since MiLAC-aided beamforming processes the symbols purely in the analog domain, each RF chain carries an individual symbol, which lies in a constellation with finite cardinality. For example, in the case of quadrature phase shift keying (QPSK), the symbols can be chosen from the constellation $\{+1+j,+1-j,-1-j,-1+j\}$, up to a scaling factor, which can be generated through DACs having just 1-bit resolution for both in-phase and quadrature parts of the signal. Thus, low-resolution ADCs/DACs can be adopted without sacrificing performance, which are less expensive and less power-hungry than high-resolution ADCs/DACs. High-resolution ADCs/DACs are instead necessary when beamforming is performed in the digital domain (entirely or in part), i.e., in digital beamforming, hybrid beamforming, and RIS-aided beamforming, as indicated in Tab. IV.
@@ V-F4 No Computations on a Per-Symbol Basis
If beamforming is performed in the digital domain (entirely or in part), the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ needs to be computed on a per-symbol basis, as shown in Tab. IV. The complexity of this operation is $8N_{RF}N_{S}$ real operations, as discussed in the Appendix of Part I, and can become prohibitive in massive/gigantic MIMO systems [4]. MiLAC-aided beamforming solves this issue since it does not require any computation on a per-symbol basis, as the beamforming is purely analog. This reduces the beamforming computational complexity, especially when the coherence time is much longer than the symbol time [4].
@@ VI-A Performance Evaluation
To evaluate the performance of R-ZFBF, ZFBF, and MBF realized with digital and MiLAC-aided beamforming, we consider a multi-user multiple-input single-output (MISO) system with an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna receivers, as presented in Section IV-B. Here, the transmitter precodes the $N_{R}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{R}\times 1}$ intended to the $N_{R}$ receivers through R-ZFBF or ZFBF to suppress inter-stream interference, or through MBF to maximize the channel gains. With digital beamforming, the R-ZFBF transmitter is realized by computing $\mathbf{F}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$ and setting the precoding matrix $\mathbf{W}_{\text{R-ZFBF}}$ such that $\mathbf{W}_{\text{R-ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, assuming uniform power allocation to the users. For large $N_{R}$ and if the signal-to-noise ratio (SNR) is the same at all receiving antennas, the optimal $\lambda$ is $\lambda=N_{R}\sigma_{\mathbf{n}}^{2}/P_{T}$, where $P_{T}$ is the transmit power, i.e., $\mathbf{s}$ has covariance matrix $\mathbf{C}_{\mathbf{s}}=P_{T}/N_{R}\mathbf{I}_{N_{R}}$, $\sigma_{\mathbf{n}}^{2}$ is the noise power, and we assumed the channel to have unit path gain [44, Chapter 12]. Similarly, digital beamforming can realize ZFBF by computing $\mathbf{F}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$ and setting $\mathbf{W}_{\text{ZFBF}}$ as $\mathbf{W}_{\text{ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, and MBF by using $\mathbf{F}_{\text{MBF}}=\mathbf{H}^{H}$ and $\mathbf{W}_{\text{MBF}}$ such that $\mathbf{W}_{\text{MBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{MBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{MBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$. When a MiLAC is used to perform R-ZFBF, ZFBF, and MBF, we have two choices. First, the MiLAC can be reconfigured to implement an arbitrary beamforming matrix, as discussed in Section IV-A. Thus, its tunable admittance components can be set according to (20) and (21), where $\mathbf{W}$ is the desired beamforming matrix $\mathbf{W}_{\text{R-ZFBF}}$, $\mathbf{W}_{\text{ZFBF}}$, or $\mathbf{W}_{\text{MBF}}$ with uniform power allocation, which is digitally computed. Second, the MiLAC can directly compute the precoding matrices of these LMMSE-inspired beamforming techniques in the analog domain, as explained in Section IV-B. In this case, the columns of $\mathbf{W}_{\text{R-ZFBF}}$, $\mathbf{W}_{\text{ZFBF}}$, and $\mathbf{W}_{\text{MBF}}$ cannot be individually normalized, and the resulting precoding matrices are given by $\mathbf{W}_{\text{R-ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{R-ZFBF}}/\|\mathbf{F}_{\text{R-ZFBF}}\|_{F}$, $\mathbf{W}_{\text{ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{ZFBF}}/\|\mathbf{F}_{\text{ZFBF}}\|_{F}$, and $\mathbf{W}_{\text{MBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{MBF}}/\|\mathbf{F}_{\text{MBF}}\|_{F}$, where the normalization factor has been included to ensure that the transmit power is the same as with digital beamforming.
@@
In Fig. 5, we report the sum rate versus the SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$ achieved by digital and MiLAC-aided beamforming, where $N_{T}=N_{R}=4$ and the channels are independent and identically distributed (i.i.d.) Rayleigh distributed with unit path gain. We observe that R-ZFBF is the best precoder, which boils down to ZFBF at high SNR and to MBF at low SNR, for both digital and MiLAC-aided beamforming, as expected from MIMO theory [44, Chapter 12]. MiLAC-aided beamforming performs exactly as digital beamforming when the MiLAC is reconfigured based on the arbitrary beamforming matrix computed digitally, denoted as MiLAC (arbit.) in Fig. 5. Besides, MiLAC-aided beamforming suffers a slight performance degradation at high SNR when the LMMSE-inspired beamforming matrix is computed in the analog domain, referred to as MiLAC (LMMSE) in Fig. 5. This occurs because, in this case, MiLAC-aided beamforming cannot perform uniform power allocation, which is optimal at high SNR. At the cost of this slight performance degradation, computing the LMMSE-inspired beamforming matrices in the analog domain substantially reduces the required computational complexity, as it will be analyzed in the following.
@@
In Fig. 6, we report the BER versus the SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$ obtained with digital and MiLAC-aided beamforming with $N_{T}=N_{R}=4$ and QPSK modulation. Since MiLAC-aided beamforming can implement the same combining matrices as digital beamforming, the two beamforming strategies achieve exactly the same BER.
@@
Figure 8: Sum rate versus computational complexity of R-ZFBF performed in a $N_{R}\times N_{T}$ system with digital and MiLAC-aided beamforming, for three different levels of SNR. Markers in blue correspond to digital beamforming and in red to MiLAC-aided beamforming.
@@ VI-C Relationship Between Performance and Computational Complexity
VI-D Performance Evaluation with Impairments
So far, we have implicitly assumed perfect channel state information (CSI) and that the tunable admittance components of the MiLAC can be reconfigured with arbitrary precision. We now compare MiLAC-aided beamforming with digital beamforming by relaxing these two assumptions, namely in the presence of noisy CSI and discrete-value admittance components in the MiLAC.
Figure 9: Sum rate in a $4\times 4$ multi-user MISO system, with noisy CSI.
Noisy CSI is modeled by adding AWGN to the perfect CSI $\mathbf{H}$, with different levels of SNR $\rho\in\{20,10,5\}$ dB. In Fig. 9, we report the sum rate achieved by R-ZFBF implemented with MiLAC-aided beamforming and digital beamforming optimized based on noisy CSI. For MiLAC-aided beamforming, we consider that the MiLAC is reconfigured to operate LMMSE-inspired beamforming, as explained in Section IV-B. We observe that with $\rho=20$ dB the performance obtained by MiLAC-aided beamforming and digital beamforming is similar to the performance in the case of perfect CSI. Besides, the performance degrades as $\rho$ decreases. MiLAC-aided beamforming approximately performs as digital beamforming for all the values of $\rho$, with a slight performance degradation at high SNR, as observed for perfect CSI in Fig. 5.
Figure 10: Sum rate in a $4\times 4$ multi-user MISO system, where the MiLAC has discrete-valued tunable admittance components.
For practical implementation, it is convenient to design a MiLAC with tunable admittance components whose values are selected from a codebook of limited cardinality. To this end, we notice that the optimal admittance components $Y_{i,k}$ with $i\neq k$ are distributed as the entries of the channel $\mathbf{H}$ when the MiLAC is used to perform R-ZFBF, following (8) and Tab. III. Thus, $Y_{i,k}$ with $i\neq k$ have real and imaginary parts Gaussian distributed in the case of i.i.d. Rayleigh fading channels, which can be optimally quantized according to [47]. In Fig. 10, we report the sum rate achieved by R-ZFBF implemented with MiLAC-aided beamforming where the admittance components $Y_{i,k}$ with $i\neq k$ are discretized with $B$ bits of resolution, where $B\in\{8,4,2\}$. For the tunable components $Y_{i,k}$ with $i\neq k$, we quantize both their real and imaginary parts with $B/2\in\{4,2,1\}$ bits of resolution according to [47]. In contrast, for the tunable components $Y_{k,k}$, we assume that they can take arbitrary values for simplicity, as they are just $N_{T}+N_{R}$. Comparing Figs. 9 and  10, we notice that the impact of discrete-value admittance components with $B\in\{8,4,2\}$ resolution bits is similar to the impact of noisy CSI with SNR $\rho\in\{20,10,5\}$ dB, respectively. This is because the signal-to-quantization-noise ratio (SQNR) obtained by optimally quantizing a real Gaussian with $B/2\in\{4,2,1\}$ bits is $\rho^{\prime}\in\{20.2,9.30,4.40\}$ dB [47], respectively, which is approximately $\rho\in\{20,10,5\}$ dB.